# Decision-Focused Learning: Foundations, State of the Art, Benchmark and Future Opportunities

### v3

@@ 2.1 Problem Setting
The goal of the CO problem above is to find a solution $\mathbf{x}^{\star}(\mathbf{c})\in\mathbb{R}^{n}$, which minimizes the objective function $f$, subject to equality and inequality constraints as defined by the functions $\bm{h}$ and $\bm{g}$. This parametric problem formulation defines $\mathbf{x}^{\star}(\mathbf{c})$ as a function of the parameters $\mathbf{c}\in\mathbb{R}^{k}$.
@@ 3.1.1 Analytical Differentiation of Optimization Mappings
$\frac{df}{d\mathbf{x}}(\mathbf{c},V(\mathbf{c}))=0$ (10)
@@
$\mathbf{x}^{\star}=\operatorname*{argmax}_{\mathbf{x}}$ $\;\;f(\mathbf{x})$ (12a)
@@
When $\mathbf{x}\in\mathbb{R}^{k}$ and the number of inequality and equality constraints are $M_{in}$ and $M_{eq}$, respectively, a QP problem is specified by parameters $Q\in\mathbb{R}^{k\times k}$, $\mathbf{c}\in\mathbb{R}^{k}$, $R\in\mathbb{R}^{k\times M_{in}}$, $\mathbf{s}\in\mathbb{R}^{M_{in}}$, $A\in\mathbb{R}^{k\times M_{eq}}$, and $\mathbf{b}\in\mathbb{R}^{M_{eq}}$. The optimal solution $\mathbf{x}^{\star}$ can be implicitly differentiated with respect to each of the parameters $(Q,\mathbf{c},R,\mathbf{s},A,\mathbf{b})$, by directly differentiating the KKT conditions of optimality (13). This produces a linear system of equations, which can be solved for the desired gradients $\frac{d\mathbf{x}^{\star}}{d\mathbf{c}}$. Later, ? (?) extended the technique of ? (?), by also computing the QP problem’s second order derivatives. This enables training with gradient boosting models, which require the gradient as well as the Hessian matrix of the loss.
@@
$\mathbf{x}^{\star}(A,\mathbf{b},\mathbf{c})=\operatorname*{argmax}_{\mathbf{x}}$ $\;\;\mathbf{c}^{\top}\mathbf{x}$ (15a)
@@ 3.1.2 Analytical Smoothing of Optimization Mappings
$\mathbf{x}^{\star}(\mathbf{c})$ $=\operatorname*{argmax}_{\mathbf{x}}\mathbf{c}^{\top}\mathbf{x}-\mu\lVert\mathbf{x}\rVert^{2}_{2}$ (19a)
$=\operatorname*{argmin}_{\mathbf{x}}\lVert\mathbf{x}-\frac{\mathbf{c}}{\mu}\rVert^{2}_{2}$ (19b)
where the above equality follows from expanding the square and cancelling constant terms, which do not affect the $\operatorname*{argmax}$. This provides an intuition as to the effect of such a quadratic regularization: it converts a LP problem into that of projecting the point $\frac{\mathbf{c}}{\mu}$ onto the feasible polytope, which results in a continuous mapping $\mathbf{\hat{c}}\to\mathbf{x}^{\star}(\mathbf{\hat{c}})$. ? (?) then train decision-focused models by solving and backpropagating the respective quadratic programming problem using the OptNet framework (?), in order to learn to predict objective parameters with minimal regret. At test time, the quadratic smoothing term is removed. ? (?) refers to such regret-based DFL with quadratically regularized linear programs as the Quadratic Programming Task Loss method (QPTL).
@@
In a related approach, ? (?) propose a general, differentiable LP solver based on log-barrier regularization. For a parametrized LP of standard form (8), gradients are computed for the regularized form in which the constraints $\mathbf{x}\geq\mathbf{0}$ are replaced with log-barrier approximations:
$\mathbf{x}^{\star}(\mathbf{c})=\operatorname*{argmin}_{\mathbf{x}}\mathbf{c}^{\top}\mathbf{x}+\lambda\sum_{i}x_{i}$ (20a)
@@
While similar in this sense to (?), this method exploits several efficiencies specific to linear programming, in which the log-barrier term serves a dual purpose of rendering (20) differentiable and also aiding its solution (?). Rather than forming and solving this regularized LP problem directly, the solver uses an interior point method to produce a sequence of log-barrier approximations to the LP’s homogenous self-dual (HSD) embedding. Early stopping is applied in the interior point method, producing a solution to (20) for some $\lambda$, which serves as a smooth surrogate problem for differentiation. A major advantage of this technique is that it only requires optimization of a linear program, making it in general more efficient than direct solution of a regularized problem as in the approaches described above.
@@ 3.3.6 Active learning algorithm for DFL
We end this section with a remark that DFL can be extended to address optimization problems beyond the ‘classical’ CO problem. For instance, ? (?) embed $K$-means clustering as a layer in a neural network by differentiating through the clustering layer, using DFL techniques. ? (?) employ DFL approach to predict the parameters of Markov decision processes (MDPs).
@@ 4 Applications of Decision-Focused Learning
In learning to rank (LTR), a machine learning model must produce rankings of documents in response to users’ search queries, in which those most relevant to a given query are placed in the highest ranking positions. In this setting, the relevance of documents to queries is often measured empirically by historical user click rates (?). In fair learning to rank (FLTR), this relevance-based matching must be performed subject to strict constraints on the relative exposure between predefined groups. Due to the difficulty of enforcing such constraints on the outputs of a machine learning model, many FLTR frameworks resort to a two-stage approach in which prediction of query-document relevance scores is learned by a typical LTR model without constraints on fairness of exposure. At test time, the predicted relevance scores inform the objective of a separate fair ranking optimization program (?). ? (?) use DFL to unify the prediction of relevance scores with the subsequent optimization of fair rankings, in an end-to-end model trained by SPO which learns to map user queries directly to the fair ranking policies that optimize user relevance. The result is a FLTR model which outperforms previous penalty-based models in terms of both user relevance and fairness, with the ability to directly control their trade-offs by modifying the fairness constraints of the optimization layer.
@@ 5.1 Problem Descriptions
All the test problems, which are selected for benchmarking, have been previously used in the DFL literature and their datasets are publicly available. All these problems are Predict-Then-Optimize problems, meaning they encompass the two stages—prediction and optimization. Table 2 provides an overview of the experimental setups associated with each test problem, including the specification of the CO problem and the type of predictive model. Next, these test problems are described in detail.
@@ 5.1.3 Warcraft shortest path problem
5.1.4 Energy-cost aware scheduling
@@ 5.1.5 Knapsack problem
5.1.6 Diverse bipartite matching
@@ Acknowledgments
This research was partly funded by the European Research Council (ERC) under the EU Horizon 2020 research and innovation programme (Grant No 101002802, CHAT-Opt) and the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101070149, project Tuples. This research is also partially supported by NSF grants 2242931, 2232054, 2007164, and NSF CAREER award 2143706. Victor Bucarey was funded by the ANID Fondecyt Iniciacion Grant no. 11220864.
@@ A Results on All Problem Instances
In the main text, boxplots of some problem instances are presented to save space in the main text. In this section of Appendix, the boxplots of all the instances are provided. Figure A1, Figure A2, Figure A3, Figure A4, Figure A5, Figure A6, and Figure A7 display the boxplots for all the instances of the grid shortest path problem, the portfolio optimization problem, the Warcraft shortest path problem, the energy-cost aware scheduling problem, the knapsack problem, the diverse bipartite matching problem, and the subset selection problem respectively.
@@ C Details about Hyperparameter Configuration
The hyperparameters for each methodology in each experiment are selected through grid search, as described in Section 5.2. For the sake of reproducibility, this section provides the lists of optimal hyperparameter combinations found by grid search and used to evaluate all the methodologies. Table A1, Table A2, Table A3, Table A4, Table A5, Table A6 present the hyperaprameter combinations for the instances of the shortest path problem on the grid, portfolio optimization problem, Warcraft shortest path problem, energy-cost aware scheduling, knapsack problem and diverse bipartite matching problem respectively.
Deg 1 2 4 6 8
PF (lr) 1. 0.05 0.05 1. 0.05
SPO (lr) 0.01 0.05 0.1 1. 1.
DBB (lr, $\lambda$) (0.1, 10.) (0.1,1.) (0.05, 1.) (0.05, 0.1) (0.1, 0.1)
I-MLE (lr, $\lambda$, $\epsilon$, $\kappa$) (0.5, 100, 1, 5) (0.5, 100, 0.1, 5) (0.5, 100.,2.,5) (0.5, 10., 0.5, 5) (0.5, 100., 2., 5)
FY (lr, $\epsilon$) (1., 5) (0.1, 0.1) (0.1, 0.1) (0.05, 5.) (0.1, 5.)
HSD (lr, $\mu$, damping) (0.1, 0.001, 1.0) (0.1, 0.1, 0.01) (0.1, $10^{-6}$, 0.1) (0.1, 0.001, $10^{-6}$) (0.1, 10., 1.)
QPTL (lr, $\mu$) (0.1, 10.) (0.5, 10.) (0.1, 1.) (0.1, 10.) (0.1, 10.)
Listwise (lr, $\tau$) (0.1, 0.1) (0.1, 0.1) (1., 0.1) (0.1, 1.) (1., 1.)
Pairwise (lr, $\Theta$ ) (1., 1.) (0.1, 0.5) (0.1, 1.) (1., 10.) (1., 10.)
Pairwise(diff) (lr) 0.1 0.1 0.5 0.1 1.
MAP (lr) 0.1 0.1 1. 1. 1.
Table A1: Optimal Hyperparameter Combination for the shortest path problems on a $5\times 5$ grid.
Deg 1 4 8 16
PF (lr) 0.01 0.05 0.1 0.05
SPO (lr) 0.5 1. 0.5 0.5
DBB (lr, $\lambda$) (1., 0.1) (1., 0.1) (1., 0.1) (1., 0.1)
I-MLE (lr, $\lambda$, $\epsilon$, $\kappa$) (0.5, 0.1, 0.1, 5) (0.5, 0.1, 0.5, 5) (0.5, 0.1, 0.05,5) (0.5, 0.1, 0.05,5)
FY (lr, $\epsilon$) (0.1, 0.01) (0.5, 0.01) (1, 0.01) (1., 2.)
QPTL (lr, $\mu$) (0.1, 10.) (0.05, 10.) (0.1, 10.) (0.05, 10.)
Listwise (lr, $\tau$) (0.1, 0.01) (0.1, 0.01) (0.1, 0.01) (0.05, 0.005)
Pairwise (lr, $\Theta$ ) (0.01, 0.01) (0.01, 0.1) (0.01, 0.01) (0.1, 0.05)
Pairwise(diff) (lr) 0.1 0.1 0.1 0.05
MAP (lr) 0.01 1. 0.05 1.
Table A2: Optimal Hyperparameter Combination for the portfolio optimization problem instances.
Image Size 30 24 18 12
PF (lr) 0.01 0.001 0.0005 0.001
SPO (lr) 0.0005 0.005 0.01 0.005
DBB (lr, $\lambda$) (0.005, 10.) (0.001, 100.) (0.001, 10.) (0.001, 10.)
I-MLE (lr, $\lambda$, $\epsilon$, $\kappa$) (.001, 100, 0.05, 50) (0.001, 10., 0.05, 50) (0.01, 10., 0.05, 5) (0.001, 10., 0.05, 50)
FY (lr, $\epsilon$) (0.001, 0.01) (0.01, 0.01) (0.01, 0.01) (0.01, 0.01)
Listwise (lr, $\tau$) (0.005, 1.) (0.005, 0.5) ( 0.005, 0.05) (0.005, 0.5)
Pairwise (lr, $\Theta$ ) (0.01, 0.1) (0.01 ,0.1) (0.005, 0.1) (0.01, 0.1)
Pairwise(diff) (lr) 0.005 0.005 0.005 0.005
MAP (lr) 0.01 0.005 0.005 0.005
Table A3: Optimal Hyperparameter Combination for the Warcraft shortest path problem instances.
Instance 1 2 3
PF (lr) 0.5 0.5 0.5
SPO (lr) 1. 0.5 0.5
DBB (lr, $\lambda$) (0.01, 0.1) (0.5, 1.) (0.5, 1.)
I-MLE (lr, $\lambda$, $\epsilon$, $\kappa$) (0.5, 1., 2., 5) (0.5, 1., 1., 5) (0.5, 1., 1., 5)
FY (lr, $\epsilon$) (0.01, 0.1) (0.5, 5) (0.01, 0.1)
HSD (lr, $\mu$, damping) (0.1, 0.1, $10^{-6}$ ) (0.1, 0.001, $10^{-6}$) (0.1, 0.1, 0.1)
QPTL (lr, $\mu$) (0.1, 1.) (0.1, 1.) (0.1, 1.)
Listwise (lr, $\tau$) (0.1, 5.) (0.1, 5.) (0.1, 5.)
Pairwise (lr, $\Theta$ ) (0.1, 1.) (0.1, 5.) (0.1, 50.)
Pairwise(diff) (lr) 0.5 0.5 0.1
MAP (lr) 0.5 0.5 0.5
Table A4: Optimal Hyperparameter Combination for the energy-cost aware scheduling problem instances.
Capacity 60 120 180
PF (lr) 0.5 1. 1.
SPO (lr) 0.5 1. 1.
DBB (lr, $\lambda$) (0.5, 0.1) (1., 1.) (0.5, 1.)
I-MLE (lr, $\lambda$, $\epsilon$, $\kappa$) (0.5, 0.1, 0.5, 5) (0.5, 0.1, 0.1, 5) (0.5, 0.1, 5., 5)
FY (lr, $\epsilon$) (1., 0.005) (1., 0.5) (0.5, 0.5)
HSD (lr, $\mu$, damping) (0.5, 0.01, 10.) (0.5, 0.1, 10.) (1., 0.01, 0.1)
QPTL (lr, $\mu$) (0.5, 10.) (0.5, 1.) (0.5, 0.1)
Listwise (lr, $\tau$) (1., 0.001) (1., 0.001) (0.5, 0.0001)
Pairwise (lr, $\Theta$ ) (0.5, 10.) (0.5, 10.) (0.5, 10.)
Pairwise(diff) (lr) 1. 1. 1.
MAP (lr) 1. 1. 1.
Table A5: Optimal Hyperparameter Combination for the knapsack problem instances.
$(\rho_{1},\rho_{2})$ (10%, 10%) (25%, 25%) (50%, 50%)
PF (lr) 0.01 0.01 0.0005
SPO (lr) 0.001 0.001 0.005
DBB (lr, $\lambda$) (0.01, 10.) (0.01, 0.1) (0.01, 1.)
I-MLE (lr, $\lambda$, $\epsilon$, $\kappa$) (0.001, 100, 0.5, 5) (0.001, 100., 0.5, 5) ( 0.001, 100., 0.5, 5)
FY (lr, $\epsilon$) ( 0.001, 0.5) (0.001, 0.01) (0.001, 5.)
HSD (lr, $\mu$, damping) (0.001, 1., 0.1) (0.05, 0.1, 10.) (0.001, 0.1, 0.1)
QPTL (lr, $\mu$) ( 0.01, 100.) (0.001. 10.) (0.001, 10.)
Listwise (lr, $\tau$) (0.001, 5.) (0.01, 5.) (0.01, 5.)
Pairwise (lr, $\Theta$ ) ( 0.005, 5) (0.01, 50.) (0.01, 50.)
Pairwise(diff) (lr) 0.001 0.01 0.005
MAP (lr) 0.001 0.01 0.005
Table A6: Optimal Hyperparameter Combination for the diverse bipartite matching problem instances.

### v4

@@ 2.1 Problem Setting
The goal of the CO problem above is to find a solution $\mathbf{x}^{\star}(\mathbf{c})\in\mathbb{R}^{n}$ ($n$ being the dimension of the decision variable $x$), which minimizes the objective function $f$, subject to equality and inequality constraints as defined by the functions $\bm{h}$ and $\bm{g}$. This parametric problem formulation defines $\mathbf{x}^{\star}(\mathbf{c})$ as a function of the parameters $\mathbf{c}\in\mathbb{R}^{m}$.
@@ 3.1.1 Analytical Differentiation of Optimization Mappings
$\frac{d}{d\mathbf{x}}f(\mathbf{c},V(\mathbf{c}))=0$ (10)
@@
$\mathbf{x}^{\star}=\operatorname*{argmin}_{\mathbf{x}}$ $\;\;f(\mathbf{x})$ (12a)
@@
When $\mathbf{x}\in\mathbb{R}^{n}$ and the number of inequality and equality constraints are $M_{in}$ and $M_{eq}$, respectively, a QP problem is specified by parameters $Q\in\mathbb{R}^{n\times n}$, $\mathbf{c}\in\mathbb{R}^{n}$, $R\in\mathbb{R}^{n\times M_{in}}$, $\mathbf{s}\in\mathbb{R}^{M_{in}}$, $A\in\mathbb{R}^{n\times M_{eq}}$, and $\mathbf{b}\in\mathbb{R}^{M_{eq}}$. The optimal solution $\mathbf{x}^{\star}$ can be implicitly differentiated with respect to each of the parameters $(Q,\mathbf{c},R,\mathbf{s},A,\mathbf{b})$, by directly differentiating the KKT conditions of optimality (13). This produces a linear system of equations, which can be solved for the desired gradients $\frac{d\mathbf{x}^{\star}}{d\mathbf{c}}$. Later, ? (?) extended the technique of ? (?), by also computing the QP problem’s second order derivatives. This enables training with gradient boosting models, which require the gradient as well as the Hessian matrix of the loss.
@@
$\mathbf{x}^{\star}(A,\mathbf{b},\mathbf{c})=\operatorname*{argmin}_{\mathbf{x}}$ $\;\;\mathbf{c}^{\top}\mathbf{x}$ (15a)
@@ 3.1.2 Analytical Smoothing of Optimization Mappings
$\mathbf{x}^{\star}(\mathbf{c})$ $=\operatorname*{argmin}_{\mathbf{x}}\mathbf{c}^{\top}\mathbf{x}+\mu\lVert\mathbf{x}\rVert^{2}_{2}$ (19a)
$=\operatorname*{argmin}_{\mathbf{x}}\lVert\mathbf{x}-\Big{(}\frac{-\mathbf{c}}{\mu}\Big{)}\rVert^{2}_{2}$ (19b)
where the above equality follows from expanding the square and cancelling constant terms, which do not affect the $\operatorname*{argmax}$. This provides an intuition as to the effect of such a quadratic regularization: it converts a LP problem into that of projecting the point $\Big{(}\frac{-\mathbf{c}}{\mu}\Big{)}$ onto the feasible polytope, which results in a continuous mapping $\mathbf{c}\to\mathbf{x}^{\star}(\mathbf{c})$. ? (?) then train models in decision-focused paradigm by solving and backpropagating the respective quadratic programming problem using the OptNet framework (?), in order to learn to predict objective parameters with minimal regret. At test time, the quadratic smoothing term is removed. ? (?) refers to such regret-based DFL with quadratically regularized linear programs as the Quadratic Programming Task Loss method (QPTL).
@@
In a related approach, ? (?) propose a general, differentiable LP solver based on log-barrier regularization. For a parametrized LP of standard form (8), gradients are computed for the regularized form in which the constraints $\mathbf{x}\geq\mathbf{0}$ are replaced with log-barrier approximations as follows:
$\mathbf{x}^{\star}(\mathbf{c})=\operatorname*{argmin}_{\mathbf{x}}\mathbf{c}^{\top}\mathbf{x}-\mu\sum_{i=1}^{\dim(\mathbf{x})}\log(x_{i})$ (20a)
@@
While this method, in this sense, is similar to the approach of ? (?), it exploits several efficiencies specific to linear programming, in which the log-barrier term serves a dual purpose of rendering (20) differentiable and also aiding its solution (?). Rather than forming and solving this regularized LP problem directly, the solver uses an interior point method to produce a sequence of log-barrier approximations to the LP’s homogenous self-dual (HSD) embedding. Early stopping is applied in the interior point method, producing a solution to (20) for some $\mu$, which serves as a smooth surrogate problem for differentiation. A major advantage of this technique is that it only requires optimization of a linear program, making it in general more efficient than direct solution of a regularized problem as in the approaches described above.
@@ 3.3.6 Active Learning Algorithm for DFL
We end this section with a remark that DFL can be extended to address optimization problems beyond the ‘classical’ CO problem, as defined in (1). For instance, ? (?) embed $K$-means clustering as a layer in a neural network by differentiating through the clustering layer, using DFL techniques. ? (?) employ a DFL approach to predict the parameters of Markov decision processes (MDPs).
@@ 4 Applications of Decision-Focused Learning
In learning to rank (LTR), a machine learning model must produce rankings of documents in response to users’ search queries, in which those most relevant to a given query are placed in the highest ranking positions. In this setting, the relevance of documents to queries is often measured empirically by historical user click rates (?). In fair learning to rank (FLTR), this relevance-based matching must be performed subject to strict constraints on the relative exposure between predefined groups. Due to the difficulty of enforcing such constraints on the outputs of a machine learning model, many FLTR frameworks resort to a two-stage approach in which prediction of query-document relevance scores is learned by a typical LTR model without constraints on fairness of exposure. At test time, the predicted relevance scores inform the objective of a separate fair ranking optimization program (?). ? (?) use DFL to unify the prediction of relevance scores with the subsequent optimization of fair rankings, in an end-to-end model trained by SPO. This DFL approach to FLTR requires significant computation to solve a linear program for fair rankings throughout training; it was subsequently proposed in ? (?) to employ a more efficient optimization of ordered weighted averages following ? (?), adapted to leverage the SPO training framework.
@@ 5.1 Problem Descriptions
All the test problems, which are selected for benchmarking, have been previously used in the DFL literature and their datasets are publicly available. All these problems are Predict-Then-Optimize problems, meaning they encompass the two stages—prediction and optimization. Moreover, all the optimization problems have linear objective functions. Table 2 provides an overview of the experimental setups associated with each test problem, including the specification of the CO problem and the type of predictive model. Next, these test problems are described in detail.
@@ 5.1.3 Warcraft Shortest Path Problem
5.1.4 Energy-Cost Aware Scheduling Problem
@@ 5.1.5 Knapsack Problem
5.1.6 Diverse Bipartite Matching Problem
@@ Acknowledgments
This research is partly funded by the European Research Council (ERC) under the EU Horizon 2020 research and innovation programme (Grant No 101002802, CHAT-Opt) and the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101070149, project Tuples. This research is also partially supported by NSF grants 2242931, 2232054, 2007164, and NSF CAREER award 2143706. Jayanta Mandi is supported by the Research Foundation-Flanders (FWO) project G0G3220N. Senne Berden is a fellow of the Research Foundation-Flanders (FWO-Vlaanderen, 11PQ024N). Victor Bucarey was funded by the ANID Fondecyt Iniciacion Grant no. 11220864.
@@ A Results on All Problem Instances
In the main text, boxplots of only selected problem instances are presented to save space in the main text. In this section of Appendix, the boxplots of all the instances are provided. Figure A1, Figure A2, Figure A3, Figure A4, Figure A5, Figure A6, and Figure A7 display the boxplots for all the instances of the grid shortest path problem, the portfolio optimization problem, the Warcraft shortest path problem, the energy-cost aware scheduling problem, the knapsack problem, the diverse bipartite matching problem, and the subset selection problem respectively.
@@ C Details about Hyperparameter Configuration
The hyperparameters for each methodology in each experiment are selected through grid search, as described in Section 5.2. For the sake of reproducibility lists of optimal hyperparameter combinations found by grid search and used to evaluate all the methodologies has been provided in the repository: https://github.com/PredOpt/predopt-benchmarks.